# Rapid Bayesian identification of sparse nonlinear dynamics from scarce and noisy data

### v1

@@ 2.2 Fast Bayesian Regression under the Gaussian assumption
In some fields, such as control engineering, the Bayesian approach is avoided because accurate calculations of the statistical distributions are analytically intractable, forcing one to resort to expensive Monte Carlo simulations or other sampling techniques. In particular, when the dictionary $\mathbf{D}(\mathbf{X})$ consists of nonlinear functions of $\mathbf{X}$, the distribution of each entry in $\mathbf{D}(\mathbf{X})$ is probably non-Gaussian. Nonetheless, if we accept the Gaussian approximation in the previous section, then the likelihood can be approximated as Gaussian. Combining the Gaussian likelihood with a Gaussian prior, one can maintain analytical tractability throughout the calculation. This is effectively how SparseBayes obtained its fast computational efficiency.
In the following section, we will introduce the Bayesian view of regression and derive our regression method alongside SparseBayes. For pedagogical reasons, in this section we will focus on the regression of a single state variable, i.e. a single column in $\mathbf{t}$ (and $\mathbf{w}$). One can easily extend the method to multiple dimensions later by repeating the regression for each state variable.
A Bayesian view of regression is as follows. In the Bayesian framework, instead of seeking $\mathbf{w}$ that best fits (12), one seeks the posterior distribution of $\mathbf{w}$ conditioned on the dataset $(\mathbf{D},\mathbf{t})$, which can be calculated using the Bayes rule
$p(\mathbf{w}|\mathbf{t},\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\frac{p(\mathbf{t}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})p(\mathbf{w}|\boldsymbol{\alpha})}{p(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})}.$ (13)
@@
$\mathcal{J}(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=-\frac{1}{2}(\ln|{2\pi}\mathbf{C}|+\mathbf{t}^{T}\mathbf{C}^{-1}\mathbf{t}).$ (14)
@@
Now, during regression, one can perform model selection by selecting the columns $\mathbf{d}_{m}$ in $\mathbf{D}$ that should be included in the model while removing columns that should not be in the model. In STLS, columns are removed by setting the corresponding $w_{m}$ to zero if it is smaller than a certain threshold. In a Bayesian framework, models are compared by the evidence, so we seek to maximise the log-evidence $\mathcal{J}(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$ by removing columns from $\mathbf{D}$ In the calculation in (15), this is equivalent to setting the corresponding elements in $\boldsymbol{\alpha}$ to infinity. Now, it is interesting to isolate how each element $\alpha_{m}$ and the corresponding column $\mathbf{d}_{m}$ changes $\mathcal{J}(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$. By defining
@@
and using the Sherman-Morrison formula, we isolate the contribution of the $m$-th term $\mathcal{J}(\alpha_{m})$ from the evidence without $m$-th term $\mathcal{J}(\boldsymbol{\alpha}_{-m})$, i.e. $\mathcal{J}(\mathbf{t}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\mathcal{J}(\boldsymbol{\alpha}_{-m})+\mathcal{J}(\alpha_{m})$, and find that
$\mathcal{J}(\alpha_{m})=\frac{1}{2}\left(\ln{\alpha_{m}}-\ln{(\alpha_{m}+\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m})+\frac{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{t})^{2}}{\alpha_{m}+\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}}}\right).$ (17)
Figure 1: Contribution of the $m$-th term in the model to the evidence at given prior variance $\alpha_{m}$. Data are generated from a Lorenz system, with noise $\sigma_{x}=0.25$ added to 125 data points. Terms in the model are highlighted by their respective colours, whereas terms that should not be in the model are grey. The inset shows the zoomed-in picture of the same graph near $\mathcal{J}(\alpha_{m})=0$. Notice that terms that should be included have dramatically positive $\mathcal{J}(\alpha_{m})$, except for $x_{2}$ in $(b)$, which has a relatively small but positive contribution. Also, in the inset of $(a)$, some grey lines are slightly above zero in small regions of $\alpha_{m}$, but remain negative when $\alpha_{m}$ is small enough. For our method, $\alpha_{m}$ is fixed at a small value, as shown by the black dashed line.
@@
$\alpha_{m}=\frac{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m})^{2}}{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{t})^{2}-\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}}.$ (18)
@@
Therefore, one can determine if a term should be included in the model by observing if $\mathcal{J}(\alpha_{m})>0$ for some $\alpha_{m}$, which can be determined by $(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{t})^{2}-\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}>0$. However, it should be noted that $\mathcal{J}(\alpha_{m})$ depends on $\mathbf{C}_{-m}$, which changes according to the existence of other terms in the model. As other columns $\mathbf{d}_{i\neq m}$ are added or subtracted from the model, $\mathcal{J}(\alpha_{m})$ for the column $\mathbf{d}_{m}$ will need to be recomputed.
@@
The SparseBayes algorirthm exploits this observation and devises an iterative algorithm that optimises each element in $\boldsymbol{\alpha}$ to maximise $\mathcal{J}(\alpha_{m})$, and removes incorrect terms that do not contribute to evidence positively, thereby promoting sparsity. Another interpretation is that this algorirthm effectively makes the Gaussian prior a sparsity-promoting spike-and-slab prior. Since the algorithm always seeks to maximise $\mathcal{J}(\alpha_{m})$ for each $\alpha_{m}$, incorrect terms will have the corresponding $\alpha_{m}$ optimised to $\alpha_{m}\rightarrow\infty$, effectively making the prior a spike at zero, while terms that should be included in the model will have a finite $\alpha_{m}$ value, which is effectively a slab prior. As for $\boldsymbol{\beta}$, the SparseBayes algorithm accepts a prescribed value $\sigma^{2}$ for all data points, i.e. $\mathbf{B}=\sigma^{-2}\mathbf{I}$. Alternatively, the algorithm can optimise a single $\sigma^{2}$ value for maximum evidence, but it does not accept $\sigma^{2}$ varying across different time points.
@@
Here, we present a novel Bayesian regression approach that also maximises evidence, but differs from SparseBayes in the optimisation of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$. Phenomenologically, we noticed that in some cases, incorrect terms have $\mathcal{J}(\alpha_{m})$ values exceeding zero in some small regions of $\alpha$, which might be attributed to incidental matching between spurious terms and noise. SparseBayes often mistakenly captures these spurious terms, usually with small coefficients. Zhang & Lin [28] addressed this problem by incorporating a thresholding procedure following the regression performed by the SparseBayes algorithm. In contrast, this study introduces a technique that circumvents the optimisation process of $\boldsymbol{\alpha}$ entirely. Instead, based on the observation that $\mathcal{J}(\alpha_{m})$ for the correct terms is always positive for a large range of $\alpha_{m}$, we set $\boldsymbol{\alpha}$ to a certain low value, ensuring that $\mathcal{J}(\alpha_{m})$ remains positive for the correct terms, while becoming negative for the incorrect ones. The value of $\boldsymbol{\alpha}$ can often be approximated by expert prior knowledge or physical scaling argument, and it represents our prior knowledge of the system. Often the value of $\boldsymbol{\alpha}$ is not critical, as long as it is small enough to ensure that $\mathcal{J}(\alpha_{m})$ remains positive for the correct terms. In another viewpoint, $\boldsymbol{\alpha}$ is effectively a tunable hyperparameter in our proposed method.
Another distinguishing feature of our method compared to SparseBayes and previous works [28, 27, 4] is its treatment of the noise inverse variance $\boldsymbol{\beta}$. Previous methods assumed that the noise level is the same across all data points, but this assumption may impede the algorithm’s ability to learn the optimal outcome. Instead, (15) showed a more precise description of the noise variance in the context of learning dynamic equations from time-series data. The formulation requires the estimated parameter $\mathbf{w}$ a priori, but its value remains unknown until after the regression. Hence, we propose an algorithm for an iterative technique to determine $\boldsymbol{\beta}$ by considering the $\mathbf{w}$-dependent component of equation (9) as a correction. Initially, we conduct regression under the assumption that $\mathbf{w}=0$, which is consistent with the prior. Following each regression, we modify the value of $\boldsymbol{\beta}=\boldsymbol{\sigma}^{-2}$ based on the value of $\mathbf{w}$ from the last iteration and equation (9). Subsequently, we iterate the regression process until the modification to $\boldsymbol{\beta}$ falls below a specific numerical threshold. Generally, convergence can be achieved within 3-5 iterations.
In order to find the optimal model, specifically the model with maximum evidence, we suggest using a Greedy algorithm. Starting from the full model comprising $M$ terms, we assess the potential increase in evidence by eliminating each individual term. Subsequently, we choose $M-1$ terms that yield the greatest increase in evidence, and iterate this procedure until any further increase in evidence is impossible. The technique of our proposed method is outlined in algorithm 1, which combines our search strategy to maximise evidence with the aforementioned noise iteration.
The computational complexity of this algorithm scales with the number of terms in the library $M$ by $\mathcal{O}(M^{2}/2)$. Although this is slower than the STLS method, it is still much faster than a brute-force search ($\mathcal{O}(2^{M})$) and any sampling technique. Furthermore, our elimination strategy is more robust than the iterative procedure employed by SparseBayes. In the SparseBayes algorithm, the search begins with a single term and additional terms are later added based on the prospective rise in evidence. Although this technique can be more computationally efficient, its inclusion of both addition and deletion of terms may result in occasional entrapment in a loop where collinear terms are repeatedly added and eliminated. Our elimination technique, in contrast, ensures convergence within $M^{2}/2$ iterations.
@@
input : Library $\mathbf{D}(\mathbf{X})$, dynamics $\mathbf{t}$, library noise variance $L_{I}\boldsymbol{\sigma}_{\Theta}^{2}$, dynamics noise variance $\boldsymbol{\sigma}_{\partial t}^{2}$, prior variance inverse $\boldsymbol{\alpha}$
@@ 4             Set $\mathbf{\tilde{D}}$ that includes all columns in $\mathbf{D}$ except the current $i$th column;
5             Perform ridge regression $\min_{\mathbf{w}}(\tfrac{1}{2}||\boldsymbol{\beta}(\mathbf{t}-\mathbf{D}\mathbf{w})||^{2}_{2}+\tfrac{1}{2}||\boldsymbol{\alpha}\mathbf{w}||^{2}_{2})$ to get $\mathbf{w}$, using $\boldsymbol{\beta}=\boldsymbol{\sigma}_{\partial t}^{-2}$ as the inverse noise variance and the prescribed $\boldsymbol{\alpha}$ as inverse prior variance;
@@ 7                   Update $\boldsymbol{\beta}$ using (9), with $\mathbf{w}$ from the previous iteration;
8                   Perform ridge regression $\min_{\mathbf{w}}(\tfrac{1}{2}||\boldsymbol{\beta}(\mathbf{t}-\mathbf{D}\mathbf{w})||^{2}_{2}+\tfrac{1}{2}||\boldsymbol{\alpha}\mathbf{w}||^{2}_{2})$ to update $\mathbf{w}$, using the updated $\boldsymbol{\beta}$ as the inverse noise variance and the prescribed $\boldsymbol{\alpha}$ as inverse prior variance;
@@
10            Calculate and store the evidence $\mathcal{J}(\mathbf{t}|\mathbf{\tilde{D}},\boldsymbol{\beta},\boldsymbol{\alpha})$ using (14);

### v2

@@ 2.2 Fast Bayesian Regression under the Gaussian assumption
In some fields, such as control engineering, the Bayesian approach is avoided because accurate calculations of the statistical distributions are analytically intractable, forcing one to resort to expensive Monte Carlo simulations or other sampling techniques. Nonetheless, if we accept the Gaussian approximation for the likelihood as discussed in (11) and combine it with a Gaussian prior, we can maintain analytical tractability throughout the rest of the calculation. Furthermore, for model discovery, a zero-mean Gaussian prior is a good proxy for an objective prior. In theory, an objective prior for the parameter $\mathbf{w}$ can be defined hierarchically, in which the prior on the parameter $\mathbf{w}$ is given by the zero-mean Gaussian $p(\mathbf{w}|\boldsymbol{\alpha})=\mathcal{N}(0,\boldsymbol{\alpha}^{-1})$, and a hyperprior on the inverse variance $\boldsymbol{\alpha}^{-1}$ is given by a log-uniform distribution. The hyperprior ensures that the overall prior contains no information on the scale of $\mathbf{w}$. However, previous work [36] has shown that such a hierarchical prior may significantly bias the maximum a posteriori estimate of the parameter $\mathbf{w}$. Instead, optimising the hyperparameter $\boldsymbol{\alpha}$ to maximise the evidence can provide a good estimate of integrated evidence while avoiding the bias introduced by the hierarchical prior. In other words, a zero-mean Gaussian prior with an optimal hyperparameter $\boldsymbol{\alpha}$ is a good approximate objective prior for sparse regression [24]. It is on this premise that SparseBayes was developed, while this study will also follow the same principle.
In the following section, we will introduce the Bayesian view of sparse regression and derive our regression method alongside SparseBayes. For pedagogical reasons, in this section we will focus on the regression of a single state variable, i.e. a single column in $\mathbf{y}$ (and $\mathbf{w}$). One can easily extend the method to multiple dimensions later by repeating the regression for each state variable.
In the Bayesian framework, instead of seeking $\mathbf{w}$ that best fits (11), one seeks the posterior distribution of $\mathbf{w}$ conditioned on the dataset $(\mathbf{D},\mathbf{y})$, which can be calculated using Bayes’ rule
$p(\mathbf{w}|\mathbf{y},\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\frac{p(\mathbf{y}|\mathbf{D},\mathbf{w},\boldsymbol{\beta})p(\mathbf{w}|\boldsymbol{\alpha})}{p(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})}.$ (12)
@@
$\mathcal{J}(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=-\frac{1}{2}(\ln|{2\pi}\mathbf{C}|+\mathbf{y}^{T}\mathbf{C}^{-1}\mathbf{y}).$ (13)
@@
Now, during regression, one can perform model selection by selecting the columns $\mathbf{d}_{m}$ in $\mathbf{D}$ that should be included in the model while removing columns that should not be in the model. In STLS, columns are removed by setting the corresponding $w_{m}$ to zero if it is smaller than a certain threshold. In a Bayesian framework, models are compared by their evidence, so we seek to maximise the log-evidence $\mathcal{J}(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$ by removing columns from $\mathbf{D}$ In the calculation in (14), this is equivalent to setting the corresponding elements in $\boldsymbol{\alpha}$ to infinity. Now, it is interesting to isolate how each element $\alpha_{m}$ and the corresponding column $\mathbf{d}_{m}$ changes $\mathcal{J}(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})$. By defining
@@
and using the Sherman-Morrison formula, we isolate the contribution of the $m$-th term $\mathcal{J}(\alpha_{m})$ from the evidence without the $m$-th term $\mathcal{J}(\boldsymbol{\alpha}_{-m})$, i.e. $\mathcal{J}(\mathbf{y}|\mathbf{D},\boldsymbol{\beta},\boldsymbol{\alpha})=\mathcal{J}(\boldsymbol{\alpha}_{-m})+\mathcal{J}(\alpha_{m})$, and find that
$\mathcal{J}(\alpha_{m})=\frac{1}{2}\left(\ln{\alpha_{m}}-\ln{(\alpha_{m}+\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m})+\frac{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{y})^{2}}{\alpha_{m}+\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}}}\right).$ (16)
Figure 1: Contribution of the $m$-th term in the model to the evidence at given prior variance $\alpha_{m}$. Data are generated from a Lorenz system, with noise $\sigma_{x}=0.27$ added to 201 data points. Terms in the model are highlighted by their respective colours, whereas terms that should not be in the model are grey. The inset shows the zoomed-in picture of the same graph near $\mathcal{J}(\alpha_{m})=0$. Notice that terms that should be included have dramatically positive $\mathcal{J}(\alpha_{m})$, except for $x_{2}$ in $(b)$, which has a relatively small but positive contribution. Also, in the inset of $(a,c)$, some grey lines are slightly above zero in small regions of $\alpha_{m}$, but remain negative when $\alpha_{m}$ is small enough. For our method, $\alpha_{m}$ is fixed at a small value, as shown by the black dashed line.
@@
$\alpha_{m}=\frac{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m})^{2}}{(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{y})^{2}-\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}}.$ (17)
@@
Therefore, one can determine if a term should be included in the model by observing if $\mathcal{J}(\alpha_{m})>0$ for some $\alpha_{m}$, which can be determined by $(\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{y})^{2}-\mathbf{d}_{m}^{T}\mathbf{C}^{-1}_{-m}\mathbf{d}_{m}>0$. However, it should be noted that $\mathcal{J}(\alpha_{m})$ depends on $\mathbf{C}_{-m}$, which changes according to the existence of other terms in the model. As other columns $\mathbf{d}_{i\neq m}$ are added or subtracted from the model, $\mathcal{J}(\alpha_{m})$ for the column $\mathbf{d}_{m}$ will need to be recomputed.
@@
The SparseBayes algorithm exploits this observation and devises an iterative algorithm that optimises each element in $\boldsymbol{\alpha}$ to maximise $\mathcal{J}(\alpha_{m})$, and removes incorrect terms that do not contribute to evidence positively, thereby promoting sparsity. Another interpretation is that this algorirthm effectively makes the Gaussian prior a sparsity-promoting spike-and-slab prior. Since the algorithm always seeks to maximise $\mathcal{J}(\alpha_{m})$ for each $\alpha_{m}$, incorrect terms will have the corresponding $\alpha_{m}$ optimised to $\alpha_{m}\rightarrow\infty$, effectively making the prior a spike at zero, while terms that should be included in the model will have a finite $\alpha_{m}$ value, which is effectively a slab prior. As for $\boldsymbol{\beta}$, the SparseBayes algorithm accepts a prescribed value $\sigma^{2}$ for all data points, i.e. $\mathbf{B}=\sigma^{-2}\mathbf{I}$. Alternatively, the algorithm can optimise a single $\sigma^{2}$ value for maximum evidence, but does not accept $\sigma^{2}$ varying across different time points.
@@
Here, we present a novel Bayesian regression approach that also maximises evidence, but differs from SparseBayes in the optimisation of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$. Phenomenologically, we noticed that in some cases, incorrect terms have $\mathcal{J}(\alpha_{m})$ values exceeding zero in some small regions of $\alpha$, which might be attributed to incidental matching between spurious terms and noise. SparseBayes often mistakenly captures these spurious terms, usually with small coefficients. Zhang & Lin [29] addressed this problem by incorporating a thresholding procedure following the regression performed by the SparseBayes algorithm. In contrast, this study introduces a technique that circumvents the optimisation process of $\boldsymbol{\alpha}$ entirely. Instead, based on the observation that $\mathcal{J}(\alpha_{m})$ for the correct terms is always positive for a large range of $\alpha_{m}$, we set $\boldsymbol{\alpha}$ to a certain low value, ensuring that $\mathcal{J}(\alpha_{m})$ remains positive for the correct terms, while becoming negative for the incorrect ones. The value of $\boldsymbol{\alpha}$ can often be approximated by expert prior knowledge or physical scaling arguments, and it represents our prior knowledge of the system. We found that often the value of $\boldsymbol{\alpha}$ is not critical, as long as it is small enough to ensure that $\mathcal{J}(\alpha_{m})$ remains positive for the correct terms. In another viewpoint, $\boldsymbol{\alpha}$ is effectively a tunable hyperparameter in our proposed method, and it represents the a weakly informative prior on the scale of the parameter $\mathbf{w}$.
@@
input : Library $\mathbf{D}(\mathbf{X})$, dynamics $\mathbf{y}$, library noise variance $L_{I}^{2}\boldsymbol{\sigma}_{\Theta}^{2}$, dynamics noise variance $L_{\partial t}^{2}\boldsymbol{\sigma}_{x}^{2}$, prior variance inverse $\boldsymbol{\alpha}$
@@ 4             Set $\mathbf{\tilde{D}}$ that includes all columns in $\mathbf{D}$ except the current $i$th column;
5             Perform ridge regression $\min_{\mathbf{w}}(\tfrac{1}{2}||\boldsymbol{\beta}(\mathbf{y}-\mathbf{D}\mathbf{w})||^{2}_{2}+\tfrac{1}{2}||\boldsymbol{\alpha}\mathbf{w}||^{2}_{2})$ to get $\mathbf{w}$, using $\boldsymbol{\beta}=(L_{\partial t}^{2}\boldsymbol{\sigma}_{x}^{2})^{-1}$ as the inverse noise variance and the prescribed $\boldsymbol{\alpha}$ as inverse prior variance;
@@ 7                   Update $\boldsymbol{\beta}$ using (8), with $\mathbf{w}$ from the previous iteration;
8                   Perform ridge regression $\min_{\mathbf{w}}(\tfrac{1}{2}||\boldsymbol{\beta}(\mathbf{y}-\mathbf{D}\mathbf{w})||^{2}_{2}+\tfrac{1}{2}||\boldsymbol{\alpha}\mathbf{w}||^{2}_{2})$ to update $\mathbf{w}$, using the updated $\boldsymbol{\beta}$ as the inverse noise variance and the prescribed $\boldsymbol{\alpha}$ as inverse prior variance;
@@
10            Calculate and store the evidence $\mathcal{J}(\mathbf{y}|\mathbf{\tilde{D}},\boldsymbol{\beta},\boldsymbol{\alpha})$ using (13);
@@
Another distinguishing feature of our method compared to SparseBayes and previous studies [29, 28, 4] is its treatment of the noise inverse variance $\boldsymbol{\beta}$. Previous methods assumed that the noise level is a constant across all data points, but this assumption may impede the algorithm’s ability to learn the optimal outcome. Instead, (8) and (14) showed a more precise, albeit not the most accurate, approximation to the noise variance, in which the variance’s dependency on the learnt parameter $\mathbf{w}$ is taken into account. The formulation requires the estimated parameter $\mathbf{w}$ a priori, but its value remains unknown until after the regression. Hence, we propose an algorithm for an iterative technique to determine $\boldsymbol{\beta}$ by considering the $\mathbf{w}$-dependent component of equation (8) as a correction. Initially, we conduct regression under the assumption that $\mathbf{w}=0$, which is consistent with the prior. Following each regression, we modify the value of $\boldsymbol{\beta}=\boldsymbol{\sigma}^{-2}$ based on the value of $\mathbf{w}$ from the last iteration and equation (8). Subsequently, we iterate the regression process until the modification to $\boldsymbol{\beta}$ falls below a specific numerical threshold. Generally, convergence can be achieved within 3-5 iterations.
In order to find the optimal model, specifically the model with maximum evidence, we suggest using a Greedy algorithm, outlined in algorirthm 1. Starting from the full model comprising $M$ terms, we assess the potential increase in evidence by eliminating each individual term. Subsequently, we choose $M-1$ terms that yield the greatest increase in evidence, and iterate this procedure until any further increase in evidence is impossible. The technique of our proposed method is outlined in algorithm 1, which combines our search strategy to maximise evidence with the aforementioned noise iteration.
The computational complexity of this algorithm scales with the number of terms in the library $M$ by $\mathcal{O}(M^{2}/2)$. Although this is slower than the STLS method, it is still much faster than a brute-force search ($\mathcal{O}(2^{M})$) and any sampling technique. Furthermore, our elimination strategy is more robust than the iterative procedure employed by SparseBayes. In the SparseBayes algorithm, the search begins with a single term and additional terms are later added based on the prospective rise in evidence. Although this technique can be more computationally efficient, its inclusion of both addition and deletion of terms may result in occasional entrapment in a loop where collinear terms are repeatedly added and eliminated. Our elimination technique, in contrast, ensures convergence within $M^{2}/2$ iterations. It should be noted, however, that the Greedy algorithm does not guarantee convergence to the global maximum, as there is always a chance that the algorithm may prematurely remove a term that should have been included in the model. Nevertheless, the likelihood of this happening is minimal because terms that were removed early often have a substantial adverse impact on the evidence (see Figure 1). Furthermore, with respect to the convergence of Greedy algorithms, the conventional STLS algorithm employed by the original SINDy framework also exhibits similar features to the proposed algorithm. This is because the thresholding procedure is also essentially a greedy algorithm. The issue of convergence to global extrema in STLS was addressed in [37], which may also shed some light on the convergence of the proposed Greedy algorithm.